# A Scalable MCMC Algorithm for Bayesian Inference on Binary Model Spaces

### 27 Mar 2026 v2

@@ Abstract
This article considers Bayesian model inference on binary model spaces. Binary model spaces are used by a large class of models that include graphical models, variable selection, mixture distributions, and decision trees. Traditional strategies in this field, such as reversible jump or birth-death MCMC algorithms, are still popular, despite suffering from a slow exploration of the model space. In this article, we propose an alternative: the Multiple Jump MCMC algorithm. The algorithm is simple, rejection-free, and remarkably fast. When applied to undirected Gaussian graphical models, it is $100$ to $200$ times faster than the state-of-the art, solving models with $500,000$ parameters in less than a minute. We provide theorems showing how accurately our algorithm targets the posterior, and we apply our framework to Gaussian graphical models, Ising models and variable selection, but note that it applies to most Bayesian posterior inference on binary model spaces.
@@ 1 Introduction
This article introduces a new MCMC algorithm for Bayesian model inference on binary model spaces. That is, we consider a set of candidate models, each with an associated parameter vector, and quantify the posterior for each model. We assume that each candidate model can be represented by a binary vector. This set-up is used by a large class of Bayesian model inference problems, including variable selection (Tadesse and Vannucci, 2021), mixture distributions (Stephens, 2000), regression trees (Mohammadi et al., 2020) and graphical models. Examples of graphical models include undirected Gaussian graphical models (Vogels et al., 2024), Ising models (Marsman et al., 2022) or mixed graphical models (Wang et al., 2023).
@@
On binary model spaces (and on general model spaces) Markov chain Monte Carlo (MCMC) methods are a popular tool for Bayesian model inference. These methods often circumvent the computation of normalizing constants and are suitable for exploring larger model spaces. MCMC methods design a Markov chain that converges to the posterior distribution. When the Markov chain is run long enough, the states in the tail of the Markov chain correspond to samples from the posterior.
@@
MCMC methods generally focus on uncovering the joint posterior, i.e. the posterior distribution of both the model and the associated parameters. When the associated parameter vectors have different dimensions, these MCMC methods are also referred as trans-dimensional. They exist in discrete-time (Green, 1995) or continuous-time (Mohammadi and Wit, 2015). Alternatively, one can integrate out the parameter vector and focus instead on the marginal posterior. This strategy creates an MCMC solely over the model space, see Madigan and York (1995) for a discrete-time example and Mohammadi et al. (2024) for a continuous-time example.
@@
Both trans-dimensional and marginal methods are either reversible or non-reversible. Reversible methods depend on a condition called detailed balance, which states that every move must balance with its reverse move. The detailed balance condition is more than is needed to prove convergence, but a convenient restriction to impose. Popular examples of reversible algorithms are the Reversible Jump MCMC (Green, 1995) and the Metropolis-Adjusted Langevin Algorithm (Besag, 1994; Schreck et al., 2013). Both techniques, however, depend on an accept/reject step in which proposed moves can be rejected, thereby wasting computational power. Birth-death (BD) algorithms are a popular reversible alternative. They operate in continuous time and explore the parameter space using repeated births and/or deaths, making them rejection-free. Recent examples of BD algorithms for Bayesian model inference include Gaussian graphical models (Mohammadi et al., 2024) and Gaussian directed acyclic graphical models (Jennings and Corcoran, 2016). Despite being rejection-free, BD methods still suffer from poor mixing, because they only allow local moves.
@@
Alternatively, non-reversible algorithms relax the detailed balance condition and employ different techniques to establish convergence to the stationary distribution. They enjoy better mixing properties and therefore explore large parameter spaces. In discrete time, a popular non-reversible MCMC strategy involves the idea of “lifting” (Sherlock and Thiery, 2021; Gagnon and Doucet, 2020). In continuous-time, popular non-reversible algorithms are the Zig-Zag algorithm (Bierkens et al., 2016) and the Bouncy Particle Sampler (Bouchard-Côté et al., 2018). Although both the Zig-Zag algorithm and the Bouncy Particle Sampler depend on differentiable parameter spaces, adaptations exists that allow for their application to binary model spaces, see for example (Bierkens et al., 2022; Chevallier et al., 2023). Despite better mixing properties, non-reversible algorithms still either require an accept/reject step or depend on local moves.
@@
The starting point of our work is a reversible BD algorithm. We show that with a simple trick, we can transform such a process into a discrete-time MCMC algorithm that we call the Multiple Jump MCMC (MJ-MCMC) algorithm. Our algorithm is non-reversible and targets the marginal posterior. Unlike traditional accept/reject approaches, our algorithm is rejection-free, and unlike BD processes, our algorithm allows global updates. In fact, the MJ-MCMC algorithm can traverse the entire model space in a single iteration. It is therefore exceptionally scalable, while maintaining high-quality inference. When applied to the field of undirected Gaussian graphical models (GGMs), our algorithm is $100$ to $200$ times faster than the state-of-the-art, solving $1000$-node problems (i.e. $2^{499500}$ potential models) in under $30$ seconds on a standard desktop. We apply our approach to undirected GGMs, Ising models, and Bayesian variable selection, but we note that the applicability of our work reaches beyond those examples. In fact, it can be applied to most Bayesian posterior inference on binary model spaces.
@@ 2 Preliminaries
Bayesian model inference aims to understand either the joint posterior (1) or the model posterior (2). In many cases this is challenging due to the normalizing constant $p(y)$. But even if the normalizing constant is available in closed form, computing the posterior for all $m\in\mathcal{M}$ is infeasible when the model space $\mathcal{M}$ is too large.
@@
A CTMC is defined by the $l\times l$ rate matrix $Q$ with elements $Q(m,m^{\prime})$ for all $m,m^{\prime}\in\mathcal{M}$. A CTMC explores the parameter space $\mathcal{M}$, each time moving from a state $m$ to a new state according to the following process. It samples, for every $m^{\prime}\not=m$, a value from an exponential distribution with rate $Q(m,m^{\prime})$. The chain then moves to the state $m^{\prime}$ that corresponds to the lowest sampled value. This value is called the waiting time and represents the time the chain spends in each state before moving to the next. The waiting time is continuous and different for each state, hence the classifier “continuous time”. Again, the detailed balance condition with respect to $\pi(m)$ is a convenient tool to design such a CTMC. It is expressed as
@@
Note the similarity between (3) and (4). Like a DTMC, we can design the rate matrix $Q$ such that detailed balance (4) holds with respect to the model posterior $p(m\mid y)$ in (2). The resulting limiting distribution $\pi(m)$ then equals $p(m\mid y)$. Running the CTMC long enough will then result in approximate samples from the posterior distribution.
@@
with
$q_{i}(m)=\min\left\{1,\frac{p(m^{i}\mid y)}{p(m\mid y)}\right\},$ (7)
@@
This BD process yields a simple but efficient algorithm. The main computational burden at each iteration is the computation of the $k$ rates $q_{i}(m)$. However, these rates can be computed in parallel. The MCMC chain will change exactly one element of the model $m$ at each iteration, thereby offering a rejection-free alternative to the Metropolis-Hastings approach. It is successfully applied to undirected GGMs (Mohammadi et al., 2024) and to directed acyclic graphs (Jennings and Corcoran, 2016). Moreover, Chen (2023)[Section 3.4] presents a framework containing the same idea and applies it to a low-scale Ising model.
@@
The first ratio on the right-hand side, $\frac{p(y\mid m^{\prime})}{p(y\mid m)}$, is called the Bayes factor. It quantifies the evidence in favor of one model $m^{\prime}$ over another model $m$. Like the marginal likelihood, the Bayes factor is rarely available in closed form; however, accurate closed-form approximations have appeared in many applications, including undirected Gaussian graphical models (Leppä-aho et al., 2017), directed Gaussian graphical models (Consonni and Rocca, 2012), Ising models (Marsman et al., 2022), and Bayesian variable selection (Chipman et al., 2001).
@@ 3 Proposed Approach
A single iteration of the BD algorithm in Lemma 2.1 is computationally feasible, because most rates are set to zero. However, the BD algorithm still struggles with large model spaces, because it updates just one element per iteration. Modern applications often treat models containing as many as one million elements, in which a BD algorithm needs millions of MCMC iterations leading to computation and memory-related issues. This is why we now present an extension of the birth-death algorithm. This extension creates a discrete-time Markov chain with superior mixing qualities. In fact, the new DTMC can cross the entire model space in a single iteration. We first present the algorithm, and show that it has two variants. Then, we provide a theorem for each variant, showing how accurately each variant targets the posterior distribution $p(m\mid y)$. We end this section with some additional remarks. Our algorithm provides samples from the marginal posterior (2), but we show that with a simple extension it can generate samples from the joint posterior (1) too.
@@ 3.1 Our algorithm
Algorithm 1 provides the pseudo-code of our proposed approach. As previously stated, Algorithm 1 takes an existing BD process as a starting point, and creates a new DTMC that explores the parameter space $\mathcal{M}$ faster. This is because it can update multiple elements in a single iteration. In fact, it can traverse the entire parameter space in a single iteration, because $P_{\varepsilon_{s}}(m,m^{\prime})>0$ provided that $\pi(m^{\prime})>0$ for any $m,m^{\prime}\in\mathcal{M}$. In each state, the size of a jump depends on the posterior probability of that state. Hence, we expect that the algorithm makes large jumps initially, but once it reaches an area with a high posterior probability, the jump size decreases, and the chain stabilizes.
@@ 2   Calculate in parallel the rates $q_{i}(m^{(s)})$ for all $i=1,\ldots,k$ using (6);
Although it is relatively straightforward to see that Algorithm 1 explores the space $\mathcal{M}$ fast, it is less obvious to assess how accurately the samples $(m^{(1)},\ldots,m^{(S)})$ target the posterior distribution $p(m\mid y)$. Section 3.2 studies the theoretical properties of Algorithm 1 concerning its accuracy. To that end, we consider two variants in terms of the sequence $\{\varepsilon_{s}\}_{s\geq 1}$. One variant uses a constant $\varepsilon_{s}=\varepsilon$, $s=1,2,\ldots$, which results in a homogeneous DTMC. A second variant is to construct $\{\varepsilon_{s}\}_{s\geq 1}$ as a decaying sequence, resulting in an inhomogeneous DTMC. Those two variants will be investigated separately.
3.2 Theoretical analysis
We investigate the behavior of the proposed algorithm. Recall that the DTMC of the algorithm is built upon the BD process with rate matrix $Q$ that has $\pi$ as its invariant distribution. In what follows, we will refer to this BD process as $(X(t))_{t\geq 0}$. We provide a theorem for both the homogeneous and inhomogeneous variant of Algorithm 1.
@@ 3.2 Theoretical analysis
In the first case, we have $\varepsilon_{s}=\varepsilon\in(0,1)$ for all $s\geq 1$. To concretize the discussion, we need some definitions. Let $\{M_{\varepsilon}^{(s)}\}_{s\geq 0}$ be the DTMC generated by Algorithm 1 with constant $\varepsilon$. Moreover, define the following quantities regarding $(X(t))_{t\geq 0}$ and $\{M_{\varepsilon}^{(s)}\}_{s\geq 1}$.
  • Define the continuous-valued waiting time $W_{m}$ as the random time the process $(X(t))_{t\geq 0}$ spends in state $m$ until it jumps to a new state,
  • Define the discrete-valued waiting time $W^{\varepsilon}_{m}$ as the random number of time steps the process $\{M_{\varepsilon}^{(s)}\}_{s\geq 1}$ spends in state $m$ until it jumps to a new state,
  • Define the jump probability $\widetilde{P}(m,m^{\prime})$ (resp. $\widetilde{P}_{\varepsilon}(m,m^{\prime})$) as the probability that the process $(X(t))_{t\geq 0}$ (resp. $\{M_{\varepsilon}^{(s)}\}_{s\geq 0}$) jumps from $m$ to $m^{\prime}$ conditional on the event that a transition occurs.
Theorem 3.1 below states several facts about the asymptotics of $\{M_{\varepsilon}^{(s)}\}_{s\geq 0}$ both in $\varepsilon$ and in time.
@@
    If the BD process is reversible, $\pi_{\varepsilon}-\pi||=O(\varepsilon)$ for any vector norm $||\cdot||$.
@@
    $\varepsilon W_{m}^{\varepsilon}\overset{\text{d}}{\to}W_{m}\quad\text{as}\quad\varepsilon\to 0,$
    that is, the waiting times of $(M_{\varepsilon}^{(s)})_{s\geq 0}$, scaled by $\varepsilon$, converge, in distribution, to the waiting times of $(X(t))_{t\geq 0}$.
@@
Theorem 3.1 (i) and (ii), show that, as $\varepsilon\to 0$, the limiting distribution of the DTMC of Algorithm 1 converges to $\pi$, the limiting distribution of the BD process. The convergence of (time-scaled) waiting times and jump probabilities, as shown in (iii) and (iv), shows that, as $\varepsilon$ gets smaller, Algorithm 1 behaves more similarly to the BD process; however, at the expense of slowing down. Selecting $\varepsilon$ is therefore a trade-off between computational efficiency and accuracy. A higher value of $\varepsilon$ directly results in better mixing, since the probability that the element $i$ flips equals $q_{i}(m)\varepsilon$. However, lower values of $\varepsilon$ result in more proximity to the original BD process. Ultimately, the choice of $\varepsilon$ depends on the desired qualities of the resulting MCMC algorithm. In Section 5.1, we compare the performance of Algorithm 1 with different values of constant $\varepsilon$ in the case of Gaussian graphical models.

### 02 Apr 2026 v3

# Multiple Jump MCMC: A Scalable Algorithm for Bayesian Inference on Binary Model Spaces

@@ Abstract
This article considers Bayesian model inference on binary model spaces. Binary model spaces are used by a large class of models, including graphical models, variable selection, mixture distributions, and decision trees. Traditional strategies in this field, such as reversible jump or birth-death MCMC algorithms, are still popular, despite suffering from a slow exploration of the model space. In this article, we propose an alternative: the Multiple Jump MCMC algorithm. The algorithm is simple, rejection-free, and remarkably fast. When applied to undirected Gaussian graphical models, it is $100$ to $200$ times faster than the state-of-the-art, solving models with $500,000$ parameters in less than a minute. We provide theorems showing how accurately our algorithm targets the posterior, and we show how to apply our framework to Gaussian graphical models, Ising models, and variable selection, but note that it applies to most Bayesian posterior inference on binary model spaces.
@@ 1 Introduction
This article introduces a new MCMC algorithm for Bayesian model inference on binary model spaces. That is, we consider a set of candidate models, each with an associated parameter vector, and quantify the posterior for each model. We assume that each candidate model can be represented by a binary vector. This set-up is used by a large class of Bayesian model inference problems, including variable selection (Tadesse and Vannucci, 2021), mixture distributions (Stephens, 2000), regression trees (Mohammadi et al., 2020), and graphical models. Examples of graphical models include undirected Gaussian graphical models (Vogels et al., 2024), Ising models (Marsman et al., 2022), or mixed graphical models (Wang et al., 2023).
@@
On binary model spaces (and on general model spaces), Markov chain Monte Carlo (MCMC) methods are a popular tool for Bayesian model inference. These methods often circumvent the computation of normalizing constants and are suitable for exploring larger model spaces. MCMC methods design a Markov chain that converges to the posterior distribution. When the Markov chain is run long enough, the states in the tail of the Markov chain correspond to samples from the posterior.
@@
MCMC methods generally focus on uncovering the joint posterior, i.e. the posterior distribution of both the model and the associated parameters. When the associated parameter vectors have different dimensions, these MCMC methods are also referred to as trans-dimensional. They exist in discrete-time (Green, 1995) or continuous-time (Mohammadi and Wit, 2015). Alternatively, one can integrate out the parameter vector and focus instead on the marginal posterior. This strategy creates an MCMC solely over the model space, see Madigan and York (1995) for a discrete-time example and Mohammadi et al. (2024) for a continuous-time example.
@@
Both trans-dimensional and marginal methods are either reversible or non-reversible. Reversible methods depend on a condition called detailed balance, which states that every move must balance with its reverse move. The detailed balance condition is more than is needed to prove convergence, but a convenient restriction to impose. Popular examples of reversible algorithms are the Reversible Jump MCMC (Green, 1995) and the Metropolis-Adjusted Langevin Algorithm (Besag, 1994; Schreck et al., 2013). Both techniques, however, depend on an accept/reject step in which proposed moves can be rejected, thereby wasting computational power. Birth-death (BD) algorithms are a popular reversible alternative. They operate in continuous time and explore the parameter space using repeated births and/or deaths, making them rejection-free. Recent examples of BD algorithms for Bayesian model inference include Gaussian graphical models (Mohammadi et al., 2024) and Gaussian directed acyclic graphical models (Jennings and Corcoran, 2016). Despite being rejection-free, BD methods still suffer from poor mixing because they only allow local moves.
@@
Alternatively, non-reversible algorithms relax the detailed balance condition and employ different techniques to establish convergence to the stationary distribution. They enjoy better mixing properties and therefore explore large parameter spaces. In discrete time, a popular non-reversible MCMC strategy involves the idea of “lifting” (Sherlock and Thiery, 2021; Gagnon and Doucet, 2020). In continuous-time, popular non-reversible algorithms are the Zig-Zag algorithm (Bierkens et al., 2016) and the Bouncy Particle Sampler (Bouchard-Côté et al., 2018). Although both the Zig-Zag algorithm and the Bouncy Particle Sampler depend on differentiable parameter spaces, adaptations exist that allow for their application to binary model spaces, see for example (Bierkens et al., 2022; Chevallier et al., 2023). Despite better mixing properties, non-reversible algorithms still either require an accept/reject step or depend on local moves.
@@
The starting point of our work is a reversible BD algorithm. We show that with a simple trick, we can transform such a process into a discrete-time MCMC algorithm that we call the Multiple Jump MCMC (MJ-MCMC) algorithm. Our algorithm is non-reversible and targets the marginal posterior. Unlike traditional accept/reject approaches, our algorithm is rejection-free, and unlike BD processes, our algorithm allows global updates. In fact, the MJ-MCMC algorithm can traverse the entire model space in a single iteration. It is therefore exceptionally scalable, while maintaining high-quality inference. When applied to the field of undirected Gaussian graphical models (GGMs), our algorithm is $100$ to $200$ times faster than the state-of-the-art, solving $1000$-node problems (i.e. $2^{499500}$ potential models) in under $30$ seconds on a standard desktop. We show how to apply our approach to undirected GGMs, Ising models, and Bayesian variable selection, but we note that the applicability of our work reaches beyond those examples. In fact, it can be applied to most Bayesian posterior inference on binary model spaces.
@@ 2 Preliminaries
Bayesian model inference aims to understand either the joint posterior (1) or the model posterior (2). In many cases, both are challenging due to the normalizing constant $p(y)$. But even if the normalizing constant is available in closed form, computing the posterior for all $m\in\mathcal{M}$ is infeasible when the model space $\mathcal{M}$ is too large.
@@
A CTMC is defined by the $l\times l$ rate matrix $Q$ with elements $Q(m,m^{\prime})$ for all $m,m^{\prime}\in\mathcal{M}$ (Anderson, 1991). A CTMC explores the parameter space $\mathcal{M}$, each time moving from a state $m$ to a new state according to the following process. It samples, for every $m^{\prime}\not=m$, a value from an exponential distribution with rate $Q(m,m^{\prime})$. The chain then moves to the state $m^{\prime}$ that corresponds to the lowest sampled value. This value is called the waiting time and represents the time the chain spends in each state before moving to the next. The waiting time is continuous and different for each state, hence the classifier “continuous time”. Again, the detailed balance condition with respect to $\pi(m)$ is a convenient tool to design such a CTMC. It is expressed as
@@
Note the similarity between (3) and (4). Like a DTMC, we can design the rate matrix $Q$ such that detailed balance (4) holds with respect to the model posterior $p(m\mid y)$ in (2). The resulting limiting distribution $\pi(m)$ then equals $p(m\mid y)$. Running the CTMC long enough will then result in samples from the posterior distribution.
@@
where
$q_{i}(m)=\min\left\{1,\frac{p(m^{i}\mid y)}{p(m\mid y)}\right\}$ (7)
@@
This BD process yields a simple and efficient algorithm. The main computational burden at each iteration is the computation of the $k$ rates $q_{i}(m)$. However, these rates can be computed in parallel. The MCMC chain will change exactly one element of the model $m$ at each iteration, thereby offering a rejection-free alternative to the Metropolis-Hastings approach. It is applied to undirected GGMs (Mohammadi et al., 2024) and to directed acyclic graphs (Jennings and Corcoran, 2016). Moreover, Chen (2023)[Section 3.4] presents a framework containing the same idea and applies it to a low-scale Ising model.
@@
The first ratio on the right-hand side, $\frac{p(y\mid m^{\prime})}{p(y\mid m)}$, is called the Bayes factor. It quantifies the evidence in favor of one model $m^{\prime}$ over another model $m$. The Bayes factor is rarely available in closed form, because the marginal likelihoods it comprises aren’t either; however, accurate closed-form approximations have appeared in many applications, including undirected Gaussian graphical models (Leppä-aho et al., 2017), directed Gaussian graphical models (Consonni and Rocca, 2012), Ising models (Marsman et al., 2022), and Bayesian variable selection (Chipman et al., 2001).
@@ 3 Proposed Approach
A single iteration of the BD algorithm in Lemma 2.1 is computationally feasible, because most rates are set to zero. However, the BD algorithm still struggles with large model spaces, because it updates just one element per iteration and is restricted to move to neighboring models only. Modern applications often treat models containing as many as one million elements, in which a BD algorithm needs millions of MCMC iterations, leading to computation and memory-related issues. This is why we now present an extension of the birth-death algorithm. This extension creates a discrete-time Markov chain with superior mixing qualities. In fact, the new DTMC can move across the entire model space in a single iteration. We first present the algorithm and show that it has two variants. Then, we provide a theorem for each variant, showing how accurately each variant targets the posterior distribution $p(m\mid y)$. We end this section with some additional remarks. Our algorithm provides samples from the marginal posterior (2), but we show that with a simple extension, it can generate samples from the joint posterior (1) too.
@@ 3.1 Our Algorithm
Algorithm 1 provides the pseudo-code of our algorithm. As previously stated, Algorithm 1 uses the rates of an existing BD process, and creates a new DTMC that explores the parameter space $\mathcal{M}$ faster. This is because it can update multiple elements in a single iteration. In fact, it can traverse the entire parameter space in a single iteration, because $P_{\varepsilon_{s}}(m,m^{\prime})>0$ provided that $\pi(m^{\prime})>0$ for any $m,m^{\prime}\in\mathcal{M}$. In each state, the size of a jump (i.e. the number of elements $m_{i}$ that are updated) depends on the posterior probability of that state. Hence, we expect that the algorithm makes large jumps initially, but once it reaches an area with a high posterior probability, the jump size decreases, and the chain stabilizes.
@@ 2   Calculate in parallel the rates $q_{i}(m^{(s)})$ for all $i=1,\ldots,k$ using (6);
Although it is relatively easy to see that Algorithm 1 needs less iterations to explore the space $\mathcal{M}$, it is less obvious to assess how accurately the samples $(m^{(1)},\ldots,m^{(S)})$ target the posterior distribution $p(m\mid y)$. Section 3.2 studies the theoretical properties of Algorithm 1 concerning its accuracy. To that end, we consider two variants in terms of the sequence $\{\varepsilon_{s}\}_{s\geq 1}$. One variant uses a constant $\varepsilon_{s}=\varepsilon$, which results in a homogeneous DTMC. A second variant uses a decaying $\{\varepsilon_{s}\}_{s\geq 1}$, resulting in an inhomogeneous DTMC.
3.2 Theoretical Support
We investigate the behavior of our MJ-MCMC algorithm. Recall that the DTMC of our algorithm is built upon the reversible BD process with rate matrix $Q$ that has $\pi$ as its invariant distribution. In what follows, we will refer to this BD process as $(X(t))_{t\geq 0}$. We provide a theorem for both the homogeneous and inhomogeneous variants of Algorithm 1.
@@ 3.2 Theoretical Support
In the first case, $\varepsilon_{s}=\varepsilon\in(0,1)$ for all $s\geq 1$. To concretize the discussion, let $\{M_{\varepsilon}^{(s)}\}_{s\geq 0}$ be the DTMC generated by Algorithm 1 with constant $\varepsilon$. Moreover, define the following quantities regarding $(X(t))_{t\geq 0}$ and $\{M_{\varepsilon}^{(s)}\}_{s\geq 1}$:
  • The continuous-valued waiting time $W_{m}$ as the random time the process $(X(t))_{t\geq 0}$ spends in state $m$ until it jumps to a new state,
  • The discrete-valued waiting time $W^{\varepsilon}_{m}$ as the random number of time steps the process $\{M_{\varepsilon}^{(s)}\}_{s\geq 1}$ spends in state $m$ until it jumps to a new state,
  • The jump probability $\widetilde{P}(m,m^{\prime})$ (resp. $\widetilde{P}_{\varepsilon}(m,m^{\prime})$) as the probability that the process $(X(t))_{t\geq 0}$ (resp. $\{M_{\varepsilon}^{(s)}\}_{s\geq 0}$) jumps from $m$ to $m^{\prime}$ conditional on the event that a transition occurs.
Theorem 3.1 below states several facts about the asymptotics of $\{M_{\varepsilon}^{(s)}\}_{s\geq 0}$ both in $\varepsilon$ and in time. Its proof is given in Supplementary Material A.2.
@@
    If the BD process is reversible, then $\pi_{\varepsilon}-\pi||=O(\varepsilon)$ for any vector norm $||\cdot||$.
@@
    $\varepsilon W_{m}^{\varepsilon}\overset{\text{d}}{\to}W_{m}\quad\text{as}\quad\varepsilon\to 0.$
    That is, the waiting times of $\{M_{\varepsilon}^{(s)}\}_{s\geq 0}$, scaled by $\varepsilon$, converge, in distribution, to the waiting times of $(X(t))_{t\geq 0}$.
@@
Theorem 3.1 (i) and (ii) show that, as $\varepsilon\to 0$, the limiting distribution of the DTMC of Algorithm 1 converges to $\pi$, the limiting distribution of the BD process. The convergence of (time-scaled) waiting times and jump probabilities, as shown in (iii) and (iv), shows that, as $\varepsilon$ gets smaller, Algorithm 1 behaves more similarly to the BD process; however, at the expense of slowing down. Selecting $\varepsilon$ is therefore a trade-off between computational efficiency and accuracy. A higher value of $\varepsilon$ directly results in a faster exploration of the model space, since the probability that the element $i$ flips equals $q_{i}(m)\varepsilon$. However, lower values of $\varepsilon$ result in more proximity to the original BD process. Ultimately, the choice of $\varepsilon$ depends on the desired qualities of the resulting MCMC algorithm. In Section 5.1, we compare the performance of Algorithm 1 with different values of constant $\varepsilon$ in the case of Gaussian graphical models.